OpenStudy (anonymous):

What is the Least Common Denominator of the fraction 4 over x plus the fraction 2 x over x minus 1 plus the fraction x plus 3 over 2 x ?

OpenStudy (anonymous):

2(x-1) ?

OpenStudy (ash2326):

\[\frac{4}{x}+\frac{2x}{x-1}+\frac{x+3}{2x}\] Write factors of each fraction's denominator \[x= x\] \[x-1=x-1\] \[2x= 2 \times x\] \[LCM= ( common\ terms ) \times ( uncommon\ terms)\] so Common terms= x (in first and third) Uncommon terms = x-1 and 2 so \[LCM= x \times (x-1) \times 2\]

OpenStudy (anonymous):

so its 2x(x-1) i dont fallow